OpenStudy (xxhomeworkhelperxx):

EASY QUESTION FOR FREE!!! MEDALS!!! FREE MEDALS! Convert the Word Problem into an Equation 1 more than 100 Equals 101

OpenStudy (anonymous):

1 + 100=101

OpenStudy (vianne):

1 + 100 = 101

OpenStudy (vianne):

Here's a website that might help you. http://www.purplemath.com/modules/translat2.htm

OpenStudy (anonymous):

1+100=101

OpenStudy (xxhomeworkhelperxx):

I already New the Answer I just wanted to give our Medals

OpenStudy (vianne):

Ahh.. So you're familiar with it correct? =) Kind of you to do that. :P